A marketing manager of a Zambian Indigenous Company took a random sample of 2000 customers who bought a newly introduced product and found that 500 customers knew about the product. After a vigorous advertisement, another sample of 1,500 of customers indicated that 700 customers knew about the new product. Test at 5% level of significance if the advertisement increased the number of customers who know about the new product


1
Expert's answer
2021-06-08T09:14:29-0400

Sample Proportion 1 "\hat{p_1}=\dfrac{500}{2000}=0.25"

Favorable Cases 1 "X_1=200"

Sample Size 1 "n_1=400"


Sample Proportion 2 "\hat{p_2}=\dfrac{700}{1500}=\dfrac{7}{15}\approx0.4667"

Favorable Cases 2 "X_2=700"

Sample Size 2 "n_2=1500"


The value of the pooled proportion is computed as

"\bar{p}=\dfrac{X_1+X_2}{n_1+n_2}=\dfrac{500+700}{2000+1500}\approx0.342857"



Significance Level "\alpha=0.05"


The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p_1\geq p_2"


"H_1: p_1<p_2"

This corresponds to a left-tailed test, and a z-test for two population proportions will be used.

(2) Rejection Region

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a left-tailed test is "z_c=-1.6449."

The rejection region for this lrft-tailed test is "R=\{z:z<-1.6449\}."


The z-statistic is computed as follows:

​

"z=\dfrac{\hat{p}_1-\hat{p}_2}{\sqrt{\bar{p}(1-\bar{p})(1/n_1+1/n_2)}}"

"\approx\dfrac{0.25-0.4667}{\sqrt{0.342857(1-0.342857)(1/2000+1/1500)}}"



"\approx-13.363891"

Since it is observed that "z=-13.363891<-1.6449=z_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "\hat{p}_1" is less than "\hat{p}_2," at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value is "p=0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "\hat{p}_1" is less than "\hat{p}_2," at the "\alpha=0.05" significance level.